40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 3 07:25:24 2011, kiwamu, Update, Green Locking, screwed up the end PDH box 
    Reply  Thu Dec 12 00:21:04 2013, Koji, Update, Green Locking, X end PDH box oscillation issue solved (Re: screwed up the end PDH box) 
Message ID: 4373     Entry time: Thu Mar 3 07:25:24 2011     Reply to this: 9455
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: screwed up the end PDH box 

 I somehow screwed up the PDH box at the X end station. 

Right now it's not working, so I am going to check and fix it today.

 

 In the last evening I found that one of the gain stages on the PDH box wasn't fully functional.

So I started investigating it and I though it was going to finish soon, but actually it wasn't so easy.

 

  The PDH box has several gain stages. So an input signal goes through a buffer, a filter, a boost and an output buffer stages sequentially.

The boost stage is supposed to have gain of 10, but I found it didn't have such gain.

In fact the gain was something like -30dB which is pretty small. Plus this boost stage was imposing an wired bump on the transfer function around 50 kHz.

I checked the voltages on some components around the boost stage and confirmed there were no strange voltage.

Then I suspected that the op-amp : LF356 had been broken for some reason. So I replaced it by LT1792 to see if it fixes the issue.

Indeed it did make it functional. However after few minutes of the replacement, it went back to the same bad condition.

I have no idea about what was going on at that time. Anyway it needs more careful investigations.

 

  I temporarily put a jumper cable on the board to skip this stage, but now the PDH lock is not healthy at all.

ELOG V3.1.3-